List the discontinuities for the function f(x) = cot(5x over 3).
A) There are no discontinuities.
B) n(pi over 2), where n is an integer
C) n(pi over 3), where n is an integer
D) n(three pi over five), where n is an integer

Answer by stanbon(75887)

List the discontinuities for the function f(x) = cot(5x over 3).
Since tan is zero at (5/3)x = zero and again at pi and 2pi,
cot(5x/3) is discontinuous at (5/3)x = 0 and at (5/3)x) = pi and (5/3)x = 2pi etc.
So discontinuites are at x = (3pi)/5 and x = (2*3pi/5) etc.
A) There are no discontinuities.
B) n(pi over 2), where n is an integer
C) n(pi over 3), where n is an integer
D) n(three pi over five), where n is an integer
--------------
Answer: D
----------------
Cheers,
Stan H.
-------------